### v4

# Architecture independent generalization bounds for overparametrized deep ReLU networks

Abstract.
We prove that overparametrized neural networks are able to generalize with a test error that is independent of the level of overparametrization, and independent of the Vapnik-Chervonenkis (VC) dimension. We prove explicit bounds that only depend on the metric geometry of the test and training sets, on the regularity properties of the activation function, and on the operator norms of the weights and norms of biases. For overparametrized deep ReLU networks with a training sample size bounded by the input space dimension, we explicitly construct zero loss minimizers without use of gradient descent, and prove a uniform generalization bound that is independent of the network architecture. We perform computational experiments of our theoretical results with MNIST, and obtain agreement with the true test error within a $22\%$ margin on average.
1. Introduction and main results
We consider a neural network with input space ${\mathcal{L}}_{0}\cong{\mathbb{R}}^{M_{0}}$, hidden layers ${\mathcal{L}}_{\ell}\cong{\mathbb{R}}^{M_{\ell}}$ with $\ell=1,\dots,L$, and output layer ${\mathcal{L}}_{L+1}\cong{\mathbb{R}}^{Q}$. Given $x^{(0)}\in{\mathcal{L}}_{0}$, we recursively define
$x^{(\ell)}=\sigma(W_{\ell}x^{(\ell-1)}+b_{\ell})$ (1.1)
with weight matrix $W_{\ell}\in{\mathbb{R}}^{M_{\ell}\times M_{\ell-1}}$ and bias vector $b_{\ell}\in{\mathbb{R}}^{M_{\ell}}$ for $\ell=1,\dots,L$. We assume that the activation function $\sigma:{\mathbb{R}}^{M}\rightarrow{\rm ran}(\sigma)\subset{\mathbb{R}}^{M}$ is Lipschitz continuous, and satisfies a growth bound
$|\sigma(x)|\leq a_{0}+a_{1}|x|$ (1.2)
for some constants $a_{0}\geq 0$, and $a_{1}>0$. In the output layer,
$x^{(L+1)}=W_{L+1}x^{(L)}+b_{L+1}$ (1.3)
we assume no activation to be present.
Let $\underline{\theta}\in{\mathbb{R}}^{K}$ denote the vector of parameters, containing the components of all weights and biases, so that
$K=\sum_{\ell=1}^{L+1}(M_{\ell}M_{\ell-1}+M_{\ell})\,.$ (1.4)
We will assume architectures with $M_{0}\geq M_{1}\geq\cdots\geq M_{L}\geq Q$. For every $\underline{\theta}\in{\mathbb{R}}^{K}$, the network defines the map
$f_{\underline{\theta}}:{\mathcal{L}}_{0}\rightarrow{\mathcal{L}}_{L+1}\;\;\;,\;\;\;x^{(0)}\mapsto x^{(L+1)}=f_{\underline{\theta}}(x^{(0)})$ (1.5)
from input to output layer.
We consider a problem in supervised learning, and assume to be given an a priori sample set
${\mathcal{S}}:=\{(x_{i}^{(0)},y_{i})\}_{i=1}^{N}$ (1.6)
where $x_{i}^{(0)}\in{\mathcal{L}}_{0}\cong{\mathbb{R}}^{M_{0}}$ and $y_{i}\in{\mathcal{L}}_{L+1}\cong{\mathbb{R}}^{Q}$, with $M_{0}\geq Q\geq 1$.
We extract two subsets of $\mathcal{S}$}, typically disjoint, one of which we refer to as the training set
${\mathcal{S}}^{train}:=\{(x_{i_{\alpha}}^{(0)},y_{i_{\alpha}})\}_{\alpha=1}^{n}\,$, (1.7)
and the other as the test set
${\mathcal{S}}^{test}:=\{(x_{i_{\beta}}^{(0)},y_{i_{\beta}})\}_{\beta=1}^{m}\,.$ (1.8)
For notational brevity, we will denote the elements of these sets by
${\mathcal{S}}^{train}=\{(x_{i}^{(0)},y_{i})\}_{i=1}^{n}$ (1.9)
and
${\mathcal{S}}^{test}=\{(\widetilde{x}_{j}^{(0)},\widetilde{y}_{j})\}_{j=1}^{m}\,$, (1.10)
upon appropriate relabeling.
Then, the network is trained by minimizing the cost (or loss) function
${\mathcal{C}}:=\frac{1}{n}\sum_{i=1}^{n}|f_{\underline{\theta}}(x_{i}^{(0)})-y_{i}|^{2}$ (1.11)
with respect to $\underline{\theta}$. Let $\underline{\theta}_{*}$ denote a global minimizer of $\mathcal{C}$} with fixed training set ${\mathcal{S}}^{train}$. Then,
${\mathcal{E}}_{\underline{\theta}_{*}}^{train}:=\frac{1}{n}\sum_{i=1}^{n}|f_{\underline{\theta}_{*}}(x_{i}^{(0)})-y_{i}|^{2}$ (1.12)
is the training error. It is zero if zero loss is attainable.
The test error is determined by evaluating ${\mathcal{S}}^{test}$ on the trained network,
${\mathcal{E}}_{\underline{\theta}_{*}}^{test}:=\frac{1}{m}\sum_{j=1}^{m}|f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})-\widetilde{y}_{j}|^{2}\,.$ (1.13)
In the work at hand, we will not assume that the probability distributions of ${\mathcal{S}}^{train}$ and ${\mathcal{S}}^{test}$ are a priori known, see the remarks in Section 2.1.
Definition 1.1.
We call a network strongly overparametrized if the training sample size is bounded by the dimension of input space, $n\leq M_{0}$.
We note that this is a reasonable assumption in contexts where good quality labeled data is difficult to acquire, such as in medical and hyperspectral imaging; see e.g. [WLC+23, SNH+24, GAL+25].
1.1. A priori bound
The difference between ${\mathcal{E}}_{\underline{\theta}_{*}}^{test}$ and ${\mathcal{E}}_{\underline{\theta}_{*}}^{train}$ is accounted for by the generalization error. We obtain the following a priori bounds in Proposition 1.2, which we strengthen significantly in Theorem 1.3. In Theorem 1.4, we further strengthen the latter to prove to be architecture independent for the specific case of overparametrized deep ReLU networks.
Given ${\mathcal{S}}^{train}$ and ${\mathcal{S}}^{test}$, the generalization error satisfies
$|{\mathcal{E}}_{\underline{\theta}_{*}}^{test}-{\mathcal{E}}_{\underline{\theta}_{*}}^{train}|={\mathcal{D}}[{\mathcal{S}}^{test},{\mathcal{S}}^{train}]\,$, (1.14)
where we refer to
${\mathcal{D}}[{\mathcal{S}}^{test},{\mathcal{S}}^{train}]:=\frac{1}{nm}\Big|\sum_{i=1}^{n}\sum_{j=1}^{m}\Big(|f_{\underline{\theta}_{*}}(x_{i}^{(0)})-y_{i}|^{2}-|f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})-\widetilde{y}_{j}|^{2}\Big)\Big|$ (1.15)
as the loss discrepancy between ${\mathcal{S}}^{test}$, ${\mathcal{S}}^{train}$. In particular,
${\mathcal{D}}[{\mathcal{S}}^{test},{\mathcal{S}}^{train}]=0\;\;\;{\text{i}f}\;\;\;{\mathcal{S}}^{test}={\mathcal{S}}^{train}\,$, (1.16)
and more generally, it vanishes whenever ${\mathcal{E}}_{\underline{\theta}_{*}}^{test}={\mathcal{E}}_{\underline{\theta}_{*}}^{train}$.
Proposition 1.2.
The loss discrepancy satisfies the a priori bound
${\mathcal{D}}[{\mathcal{S}}^{test},{\mathcal{S}}^{train}]\leq C_{0}(1+R)\;{\rm diam}({\mathcal{S}}^{test}\cup{\mathcal{S}}^{train})$ (1.17)
where $R$ is the radius of the smallest ball centered at the origin in ${\mathbb{R}}^{M_{0}}\times{\mathbb{R}}^{Q}$ containing ${\mathcal{S}}^{test}\cup{\mathcal{S}}^{train}$, and ${\rm diam}({\mathcal{S}}^{test}\cup{\mathcal{S}}^{train})$ is the diameter of ${\mathcal{S}}^{test}\cup{\mathcal{S}}^{train}$.
The constant
$C_{0}=C_{0}(a_{0},a_{1},c_{\sigma},\{\|W_{\ell}^{*}\|_{op},|b_{\ell}^{*}|\}_{\ell=1}^{L+1})$ (1.18)
depends on the operator norms of the trained weights and the Euclidean norms of the trained biases. Moreover, it depends on the parameters $a_{0},a_{1}$ controlling the growth of $\sigma$ in (1.2), and its Lipschitz constant $c_{\sigma}$.
1.1.1. Remarks
Using $b_{1}$, the origin in input space can be translated to minimize $R$. By Jung’s theorem, one then obtains $R\leq\sqrt{\frac{M_{0}}{2(M_{0}+1)}}\,{\rm diam}({\mathcal{S}}^{test}\cup{\mathcal{S}}^{train})$.
The a priori bound (1.18) on $C_{0}$ involves the product of operator norms of all weights, and can generically be very large. For a related result, see [NTS15].
1.2. Generalization bound
We can improve this result significantly in Theorem 1.3 if we bound $\mathcal{E}_{\underline{\theta}_{*}}^{test}$ separately instead of $|\mathcal{E}_{\underline{\theta}_{*}}^{test}-\mathcal{E}_{\underline{\theta}_{*}}^{train}|$, and even much further for ReLU networks in Theorem 1.4, where we prove bounds that do not depend on the depth of the network nor on the widths of its hidden layers.
Theorem 1.3.
Let $g_{\underline{\theta}_{*}}:{\mathbb{R}}^{M_{1}}\rightarrow{\mathbb{R}}^{Q}$ denote the map from the first hidden layer to the output layer for the trained network, so that
$f_{\underline{\theta}_{*}}(x)=g_{\underline{\theta}_{*}}(W_{1}^{*}x+b_{1}^{*})\,$, (1.19)
and let ${\rm Lip}(g_{\underline{\theta}_{*}})$ denote its Lipschitz constant. Let
${\mathcal{S}}^{train}_{W_{1}^{*}}$ $:=$ $\{(W_{1}^{*}x,y)\;|\;(x,y)\in{\mathcal{S}}^{train}\}$
${\mathcal{S}}^{test}_{W_{1}^{*}}$ $:=$ $\{(W_{1}^{*}\widetilde{x},\widetilde{y})\;|\;(\widetilde{x},\widetilde{y})\in{\mathcal{S}}^{test}\}\,.$ (1.20)
Then, defining
$C_{\underline{\theta}_{*}}:=\max\{1,{\rm Lip}(g_{\underline{\theta}_{*}})\}\,$, (1.21)
it follows that
${\mathcal{E}}_{\underline{\theta}_{*}}^{test}\leq C_{\underline{\theta}_{*}}^{2}\;d_{C}^{2}({\mathcal{S}}^{test}_{W_{1}^{*}}\,|\,{\mathcal{S}}^{train}_{W_{1}^{*}})\,$ (1.22)
when $\mathcal{E}_{\underline{\theta}_{*}}^{train}=0$, and in general
${\mathcal{E}}_{\underline{\theta}_{*}}^{test}\leq(1+\delta)C_{\underline{\theta}_{*}}^{2}d_{C}^{2}({\mathcal{S}}^{test}_{W_{1}^{*}}\,|\,{\mathcal{S}}^{train}_{W_{1}^{*}})+\left(1+\frac{1}{\delta}\right)\mathcal{E}_{\underline{\theta}_{*}}^{testNN}$ (1.23)
for any $\delta>0$, where
$\mathcal{E}_{\underline{\theta}_{*}}^{testNN}=\frac{1}{m}\sum_{j=1}^{m}r(x_{i(j)}^{(0)},y_{i(j)})^{2}$
is the average training error for points in $\mathcal{S}^{train}$ which are nearest neighbors to points in $\mathcal{S}^{test}$. These bounds hold with respect to the unidirectional Chamfer pseudodistance for point clouds,
$d_{C}({\mathcal{S}}^{test}_{W_{1}^{*}}\,|\,{\mathcal{S}}^{train}_{W_{1}^{*}}):=\left(\frac{1}{m}\sum_{(\widetilde{x},\widetilde{y})\in{\mathcal{S}}^{test}}\min_{(x,y)\in{\mathcal{S}}^{train}}\big(|W_{1}^{*}(\widetilde{x}-x)|+|\widetilde{y}-y|\big)^{2}\right)^{\frac{1}{2}}\,.$ (1.24)
For strongly overparametrized deep ReLU networks, zero training loss is attainable with explicitly constructible minimizers, and $C_{\underline{\theta}_{*}}$ in (1.22) is independent of the network architecture, see Theorem 1.4. In Section 3, we show empirically for randomly initialized ReLU networks trained on the MNIST data set that, for very small $n$, the bound (1.22) does not change significantly with depth.
1.2.1. Remarks
The unidirectional Chamfer pseudodistance does not satisfy the axioms of a metric. Along with its symmetrized version, it is widely used in computer vision, [HNSS+24]. It measures the approximability of ${\mathcal{S}}^{test}_{W_{1}^{*}}$ by subsets of ${\mathcal{S}}^{train}_{W_{1}^{*}}$, but not vice versa. In particular, $d_{C}({\mathcal{S}}^{test}_{W_{1}^{*}}\,|\,{\mathcal{S}}^{train}_{W_{1}^{*}})=0$ implies that ${\mathcal{S}}^{test}_{W_{1}^{*}}\subseteq{\mathcal{S}}^{train}_{W_{1}^{*}}$.
For Theorem 1.3, it is not necessary for ${\mathcal{S}}^{test}_{W_{1}^{*}}$ and ${\mathcal{S}}^{train}_{W_{1}^{*}}$ to be identically distributed. We also note that in (1.24), we have that
$W_{1}^{*}(\widetilde{x}-x)=W_{1}^{*}(P_{X^{(0)}}\widetilde{x}-x)\;\;\;{\rm and}\;\;\;P_{X^{(0)}}x=x$ (1.25)
for all $\widetilde{x}\in{\mathcal{S}}^{test}$ and $x\in{\mathcal{S}}^{train}$, where $X^{(0)}$ is a matrix with columns formed by the $x^{(0)}$ vectors in $\mathcal{S}^{train}$, and $P_{X^{(0)}}=X^{(0)}(X^{(0)})^{+}$. In other words, only the component of $x$ in the linear span of the training data are relevant for the generalization error. Moreover, we point out that for all training data in the set
${\mathcal{S}}^{train}_{W_{1}^{*}}=\{(W_{1}^{*}x,y)|(x,y)\in{\mathcal{S}}^{train}\}$ (1.26)
we have $W_{1}^{*}x=(W_{1}^{*}X^{(0)})(X^{(0)})^{+}x$ with $W_{1}^{*}X^{(0)}=Y-B_{1}^{*}$, and the vectors $(X^{(0)})^{+}x$ are coordinate unit vectors of the form $e_{\ell}=(0,\dots,0,1,0,\dots,0)$ ($\ell$-th entry is 1).
1.3. Uniform generalization bounds for zero loss ReLU
In Theorem 1.4, below, we determine an explicit bound on the constant $C_{0}$ for the important special case of strongly overparametrized deep ReLU networks. We present the construction of zero training loss minimizers without invoking gradient descent, and prove that the product of operator norms of weights in $C_{0}$ collapses, allowing for a uniform bound independent of the network architecture.
For notational convenience, we define the matrices associated to ${\mathcal{S}}^{train}$,
$X^{(\ell)}:=[x_{1}^{(\ell)}\cdots x_{n}^{(\ell)}]\;\;\;\in\;{\mathbb{R}}^{M_{\ell}\times n}$ (1.27)
and
$B_{\ell}:=[b_{\ell}\cdots b_{\ell}]\;\;\;\in\;{\mathbb{R}}^{M_{\ell}\times n}$ (1.28)
for $\ell=1\cdots,L+1$, with $M_{L+1}=Q$. Moreover,
$Y:=[y_{1}\cdots y_{n}]\;\;\;\in\;{\mathbb{R}}^{Q\times n}$ (1.29)
is the matrix of reference output vectors $y_{i}\in{\mathcal{L}}_{L+1}\cong{\mathbb{R}}^{Q}$ in output space.
We obtain the following bound for strongly overparametrized ReLU networks with non-increasing layer dimensions, yielding a uniform generalization bound which only depends on the training input data and labels. In particular, it is independent of the layer dimensions and of the number of layers.
Theorem 1.4.
Consider a strongly overparametrized network with $M_{0}\geq M_{1}\geq\cdots\geq M_{L}\geq Q$ and training sample size $n\leq M_{0}$. Assume that the matrix of training inputs $X_{0}\equiv X^{(0)}$ has full rank $n$, and that $\sigma$ is ReLU. Then, zero loss training is attainable with explicitly constructible minimizer.
The zero training loss minimizer is explicitly given by
$W_{1}^{*}$ $=$ $\left[\begin{array}[]{c}(Y-B_{L+1}^{*})(X^{(0)})^{+}\\ 0_{(M_{1}-Q)\times M_{0}}\end{array}\right]$ (1.32)
$W_{\ell}^{*}$ $=$ $[{\bf 1}_{M_{\ell}\times M_{\ell}}\;\;0_{(M_{\ell-1}-M_{\ell})\times M_{L}}]\;\;\;,\;\;\ell=2,\dots,L$
$W_{L+1}^{*}$ $=$ $[{\bf 1}_{Q\times Q}\;\;0_{Q\times(M_{L}-Q)}]$
$b_{\ell}^{*}$ $=$ $0\;\;\;,\;\;\ell=1,\dots,L$
$b_{L+1}^{*}$ $=$ $-\alpha(1,\dots,1)^{T}\;\;\;,\;\;\alpha=\max_{i,j}|(Y_{ij})_{-}|\,.$ (1.33)
Therefore, $\|W_{\ell}^{*}\|_{op}=1$ for all $\ell=2,\dots,L$, and hence, the constants in (1.21) have the values ${\rm Lip}(g_{\underline{\theta}_{*}})=1$ and $C_{\underline{\theta}_{*}}=1$.
Consequently, the generalization bound
${\mathcal{E}}_{\underline{\theta}_{*}}^{test}$ $<$ $\frac{1}{m}\sum_{(\widetilde{x},\widetilde{y})\in{\mathcal{S}}^{test}}\min_{(x,y)\in{\mathcal{S}}^{train}}\big(|W_{1}^{*}(\widetilde{x}-x)|+|\widetilde{y}-y|\big)^{2}$ (1.34)
$=$ $\;d_{C}^{2}({\mathcal{S}}^{train}_{W_{1}^{*}}\,|\,{\mathcal{S}}^{test}_{W_{1}^{*}})\,.$
follows from (1.22).
1.3.1. Remarks
Here, $\alpha$ is the modulus of the most negative coordinate among all labels $y_{j}$. We remark that in various classification problems, $y_{i}$ are chosen as unit vectors, with all components non-negative, whereby $\alpha=0$.
With straightforward modifications, we expect that a similar result can be proven for the same architecture, but for an activation given by a diffeomorphism
$\sigma:{\mathbb{R}}^{M}\rightarrow{\mathbb{R}}^{M}_{+}$ (1.35)
(for instance, a suitable mollification of ReLU), as studied in [CM].
Notably, the zero loss minimizers are not unique, see also [CE24, CM, Coo21, LDS22, ZZH+21].
2. Comparison with probabilistic bounds
In this section, we compare our results with standard approaches developed in Statistical Learning Theory.
2.1. Probabilistic setting
If it is assumed that the probability distributions for the training set ${\mathcal{S}}^{train}$ and test set ${\mathcal{S}}^{test}$ are known, and determined by probability measures $\mu_{train}$ and $\mu_{test}$ on ${\mathbb{R}}^{M_{0}}\times{\mathbb{R}}^{Q}$, respectively, then
${\mathcal{E}}_{\underline{\theta}_{*}}^{train}$ $=$ $\int d\mu_{train}(x,y)|f_{\underline{\theta}_{*}}(x)-y|^{2}$
${\mathcal{E}}_{\underline{\theta}_{*}}^{test}$ $=$ $\int d\mu_{test}(\widetilde{x},\widetilde{y})|f_{\underline{\theta}_{*}}(\widetilde{x})-\widetilde{y}|^{2}\,.$ (2.1)
The generalization error is usually analyzed by controlling the probability of the event that their difference is small. Probabilistic bounds on the generalization error in this context are studied extensively in the literature, see for instance [BFT17, NK19, NBS18, ZBH+21] and the references therein. We point out two crucial difficulties with this setting:
  1. (1)
    In practice, the probability distributions for the test and training sets are typically not a priori known.
  2. (2)
    It is usually assumed that ${\mathcal{S}}^{train}$ and ${\mathcal{S}}^{test}$ are sampled from the same set, and distributed identically, which suggests proximity of ${\mathcal{E}}_{\underline{\theta}_{*}}^{train}$ and ${\mathcal{E}}_{\underline{\theta}_{*}}^{test}$.
Point (2) introduces an extra constraint. In practice, $\mu_{train}$ and $\mu_{test}$ might not be identically distributed. For example, a network could be trained to distinguish apples from oranges with a training set ${\mathcal{S}}^{train}$ from country A. It should then be expected to have the ability to distinguish apples and oranges in a set ${\mathcal{S}}^{test}$ originating from country B that is not identically distributed to ${\mathcal{S}}^{train}$.
Because of point (1), concrete implementations commonly employ a similar, practicable version of generalization error as in Proposition 1.2. We point out that Proposition 1.2 holds for arbitrary but fixed ${\mathcal{S}}^{train}$ and ${\mathcal{S}}^{test}$, without any assumptions on their probability distributions. In Theorem 1.4, we will prove for strongly overparametrized ReLU networks that zero loss is attainable and that the generalization error then is independent of the network architecture, only depending on how well ${\mathcal{S}}^{test}$ can be approximated with subsets of ${\mathcal{S}}^{train}$ of equal cardinality.
2.2. Bounds based on the Vapnik-Chervonenkis dimension
The bound on the generalization error most widely used in the literature is an expression involving the Vapnik-Chervonenkis (VC) dimension $d_{VC}$ of the network. The statement is that with probability at least $1-\delta$,
${\mathcal{E}}_{\underline{\theta}_{*}}^{test}\leq{\mathcal{E}}_{\underline{\theta}_{*}}^{train}+O\Big(\sqrt{\frac{d_{VC}}{n}}+\sqrt{\frac{\log(1/\delta)}{n}}\Big)$ (2.2)
holds for identically distributed ${\mathcal{S}}^{train}$ and ${\mathcal{S}}^{test}$. $d_{VC}$ accounts for the maximum size of a point set allowing for all of its subsets to be distinguishable by the network.
2.2.1. Overparametrized networks
For neural networks, $d_{VC}$ has been determined in [BM03]. In particular, $d_{VC}$ typically increases with the level of overparametrization of the problem; that is, with the amount by which $K$ exceeds the overall output dimension $nQ$.
The bound (2.2) deteriorates with the increase of overparametrization and thus of $d_{VC}$; however, computational evidence shows that the generalization error remains bounded. This conundrum is a key open problem in Machine Learning, see the discussion of the double descent phenomenon in [Bel21, BHMM19]. The error term in (2.2) decreases as the number of training data $n$ increases, but this in turn lowers the level of overparametrization; in particular, in the limit $n\rightarrow\infty$, the problem becomes underparametrized. Therefore, the bound (2.2) is not effective in the overparametrized regime where control of the generalization error is most desired.
In this paper, we offer the following explanation to this phenomenon: The VC dimension describes the capacity of the network to accommodate any given set of input data. However, the bound does not account for the specific data determining the problem, and is independent of both ${\mathcal{S}}^{test}$ and ${\mathcal{S}}^{train}$. In Theorem 1.4 of this paper, we verify that the true generalization error for strongly overparametrized deep ReLU networks is determined by metric properties of the data distribution independent of the parameter count, and of the network architecture. For a discussion related to ours, we refer to [NK19].
Some intuition about the situation at hand can perhaps be gained from the following analogy: 1 liter of fluid can be filled into a 1 liter bottle, a 2 liter bottle, or a 50 liter bottle; however, in all cases, the amount of fluid remains the same. The size of the bottle corresponds to the VC dimension of the network, it grows as the overparametrization level increases. The amount of fluid corresponds to the test and training data sets, which are independent of the network architecture.
2.2.2. Zero loss in underparametrized regime
For generic training data, underparametrized networks are not able to admit zero loss. This is because the map from parameter space to output space (for given training data) is an embedding, and the zero loss minimizer is generically not contained in its range. However, training with zero loss in underparametrized networks is possible if ${\mathcal{S}}^{train}$ is sufficiently clustered, and therefore not generic, [CE25b, CM].
In [CE23, CE25a], zero loss minimizers are explicitly constructed for underparametrized deep ReLU networks with sufficiently clustered training data. It is assumed that there are $Q$ disjoint clusters corresponding to $Q$ equivalence classes, and the depth of the network is required to be at least $Q$. The weights and biases of the trained network are explicitly determined in [CE23, CE25a], without invoking gradient descent. Each layer acts by collapsing one of the clusters of training data to a point, and in the output layer, the resulting $Q$ disjoint points are matched to the vectors $y_{i}$, $i=1,\dots,Q$ that label the classes. Thereby, zero loss training is achieved.
A crucial question arises: The generalization error in this network is controlled by the bound (2.2), based on its VC dimension. The training error is zero, hence
${\mathcal{E}}_{\underline{\theta}_{*}}^{test}\leq O\Big(\sqrt{\frac{d_{VC}}{n}}+\sqrt{\frac{\log(1/\delta)}{n}}\Big)\,.$ (2.3)
Therefore, taking the limit $n\rightarrow\infty$ of infinite training sample size, (2.3) predicts that the test error must be zero, too.
However, with the construction presented in [CE23, CE25a], it is easy to create full measure sets of test data in the input layer (sufficiently separated from the training clusters) that do not produce zero loss. This might seem like a contradiction.
However, there is no contradiction; the resolution of this conundrum is that for (2.2) to hold, it is required that the test sample set ${\mathcal{S}}^{test}$ and training data set ${\mathcal{S}}^{train}$ are identically distributed. Therefore, to apply (2.3), test samples are only admissible if they are clustered with identical distribution as the training data; but then, the trained network maps them to output vectors with zero loss, as well.
To summarize, the discussion in Section 2.2.1 suggests that the effectiveness of (2.2) in the overparametrized regime is limited. On the other hand, (2.2) is effective in the underparametrized regime, where identical distribution of test and sample data is a necessary condition. If the latter is not satisfied, one can construct examples violating (2.2).
3. Experiments
To verify our theoretical bound (1.22) in Theorem 1.3, we trained shallow feedforward neural networks with layer dimensions $(784,784,10)$, no bias vectors, and ReLU activation to classify MNIST digits in three different ways: First, we constructed the zero loss minimizers in Theorem 1.4, abbreviated ZL below. We used these constructive minimizers as the initialization to train new networks (TFZL, “trained from zero loss”), to ascertain that they are stable under gradient descent. Finally, we trained networks from random initialization (RI).
All training was done using stochastic gradient descent with 70 epochs and learning rate $0.1$. The training samples are picked randomly out of MNIST, under the condition that all digits $\{0,1,\dots,9\}$ are represented; samples violating this requirement are rejected. The code is made available on the website https://github.com/patriciaewald/BCCEM.
In Figures 1 and 2 we have plotted the resulting test error $\mathcal{E}^{test}$ and the following upper bound on (1.22)
$\text{Bound }=\max\{1,\Pi_{\ell=2}^{L+1}||W_{\ell}||_{op}\}^{2}d_{C}^{2}(\mathcal{S}^{test}_{W_{1}}|\mathcal{S}_{W_{1}}^{train}),$ (3.1)
for a ReLU network with $L$ hidden layers, as a function of the size of the training data set. For the shallow neural networks, the product of norms simplifies to $W_{2}||_{op}$. For each value of $n$, there is one instance of ZL and TFZL, whereas the plotted values for RI are averaged over several random initialiations. The test set was kept fixed, with $m=20$ samples. Note that $n$ was capped at $550$ to stay within the assumptions of Theorem 1.4, whereby the data matrix $X_{0}$ needs to be full-rank.
Figure 1. Each plot shows the test error $\mathcal{E}^{test}$ and the bound (3.1) for the constructed or trained shallow networks. The plot for TFZL is very similar to the one for ZL (left).
Notably, our theoretical generalization bound yields good agreement with the constructed ZL network in the strongly overparametrized regime where network complexity based bounds are not effective [Bel21]. While the agreement with the randomly initialized networks is not as good, the bound for ZL represents an upper bound on the best generalization error:
$\min_{\underline{\theta}}\mathcal{E}_{\underline{\theta}}^{test}\leq\text{Bound for ZL}.$ (3.2)
This is particularly valuable in the data scarcity context: In Figure 2 we observe that as $n$ gets smaller (or as the system gets more overparametrized), the average test error for the randomly initialized networks gets closer to, and indeed surpasses, the test error for the explicitly constructed network (ZL).
Figure 2. This plot shows the resulting test error for a fixed test set and several shallow networks trained with training sets of different sizes $n$.
Min Max Average
ZL 1.15 1.42 1.22
TFZL 1.15 1.43 1.22
RI 29.21 48.09 34.51
Table 1. For each network, we computed the ratio $\gamma=\frac{Bound}{\mathcal{E}^{test}}$. In this table, we report the minimum, maximum, and average values of $\gamma$ for each type of trained network, in the range $n\in[20,200]$.
[…]
Figure 3. The (average) bound (3.1) computed for several different randomly initialized neural networks, with different architectures and for varying number $n$ of training samples. The missing data points for certain architectures indicate that the trained networks could not achieve $\mathcal{E}^{train}<0.001$.
3.1. Interpretation of results
We add the following remarks and observations on the experimental results:
  • The MNIST data set of handwritten digits contains a large number of pixels which are identical (white background) for all samples. This implies that the data set is contained in a linear subspace of ${\mathbb{R}}^{784}$ dimension smaller than 784. The rank of the data matrix $X_{0}$ increases with the number of training samples $n$ until $n\approx 550$, and stabilizes around that value as $n$ further increases. This implies that the MNIST data set is contained in a subspace of dimension $\approx 550$.
  • As our numerical experiments show, the theoretical zero loss bound deteriorates once $n>200$. This indicates that the bulk of the MNIST data set is concentrated on a subspace of dimension $\approx 200$ or less, while the excess dimensions $\approx 350=550-200$ account for random fluctuations. As a consequence, we estimate the ”effective dimension” of input space to be only $\approx 200$ or less, which means that the training sample set should not exceed $\leq 20$ per digit (for 10 digits) to ensure that we are in the strongly overparametrized regime.
  • For the reasons explained above, we infer that the regime $n<200$ is strongly overparametrized with high probability, and our theoretical bound is most effective here. We emphasize that this is precisely the regime in which the VC dimension based bound (2.2) fails to be effective; moreover, our bound does not require the training and test distributions to be identical. As $n>200$ increases, the strongly overparametrized regime is abandoned, and the theoretical bound (which assumes strong overparametrization) deteriorates, for both shallow and deep networks.
  • The network TFZL is obtained by initializing gradient descent with the theoretical zero loss minimizer from Theorem 1.4. We perform this calculation to provide numerical evidence that the latter is stable (is a fixed point) under gradient descent.
  • We note that while our theoretical bounds (1.34) and especially (3.1) may not be sharp, the range of values of the relative error $\gamma$ displayed in Table 1 indicates that, with an average error margin of 22%, their performance is quite strong in the strongly overparametrized regime, for the class of zero loss minima constructed in Theorem 1.4. They in particular show that the generalization error is bounded and independent of the network architecture.
  • The trained network RI obtained from random initialization produces (approximate) loss minimizers $\underline{\theta}_{*}^{RI}$ that can differ significantly from the theoretical zero loss minimizer $\underline{\theta}_{*}^{ZL}$ corresponding to (1.34) in Theorem 1.4, as the set of zero loss minima is degenerate. As shown in Table 1, the theoretical bound using $\underline{\theta}_{*}^{RI}$ instead of $\underline{\theta}_{*}^{RI}$ performs poorly because the bound (3.1) on ${\rm Lip}(g_{\underline{\theta}_{*}^{RI}})$ is suboptimal, while it is optimal for ${\rm Lip}(g_{\underline{\theta}_{*}^{ZL}})$. Improvements on the theoretical bounds on ${\rm Lip}(g_{\underline{\theta}_{*}}^{RI})$ will be addressed in future work.
Appendix A Proof of Proposition 1.2
To begin with, we have
${\mathcal{E}}_{\underline{\theta}_{*}}^{train}-{\mathcal{E}}_{\underline{\theta}_{*}}^{test}$ $=$ $\frac{1}{n}\sum_{i=1}^{n}|f_{\underline{\theta}_{*}}(x_{i}^{(0)})-y_{i}|^{2}-\frac{1}{m}\sum_{j=1}^{m}|f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})-\widetilde{y}_{j}|^{2}$ (A.1)
$=$ $\frac{1}{n}\sum_{i=1}^{n}\Big(|f_{\underline{\theta}_{*}}(x_{i}^{(0)})-y_{i}|^{2}-\frac{1}{m}\sum_{j=1}^{m}|f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})-\widetilde{y}_{j}|^{2}\Big)$
$=$ $\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}\Big(|f_{\underline{\theta}_{*}}(x_{i}^{(0)})-y_{i}|^{2}-|f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})-\widetilde{y}_{j}|^{2}\Big)$
where we used $1=\frac{1}{n}\sum_{i=1}^{n}1$ to pass to the second line, and $1=\frac{1}{m}\sum_{i=1}^{m}1$ to pass to the third line. Thus, (1.14) follows.
Next, we note that
$\Big|\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}\Big(|f_{\underline{\theta}_{*}}(x_{i}^{(0)})-y_{i}|^{2}-|f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})-\widetilde{y}_{j}|^{2}\Big)\Big|$ (A.2)
$=$ $\Big|\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}\Big((f_{\underline{\theta}_{*}}(x_{i}^{(0)})-f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)}))+(\widetilde{y}_{j}-y_{i})\Big)$
$\hskip 56.9055pt\cdot\Big((f_{\underline{\theta}_{*}}(x_{i}^{(0)})+f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)}))-(\widetilde{y}_{j}+y_{i})\Big)\Big|$
$\leq$ $\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}\Big(|f_{\underline{\theta}_{*}}(x_{i}^{(0)})-f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})|+|\widetilde{y}_{j}-y_{i}|\Big)$
$\hskip 56.9055pt\Big(|f_{\underline{\theta}_{*}}(x_{i}^{(0)})+f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})|+|\widetilde{y}_{j}+y_{i}|\Big)$
Using $\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}1=1$, we obtain the upper bound
$\leq$ $\sup_{(x,y)\in{\mathcal{S}}^{train}}\sup_{(\widetilde{x},\widetilde{y})\in{\mathcal{S}}^{test}}\Big(|f_{\underline{\theta}_{*}}(x)-f_{\underline{\theta}_{*}}(\widetilde{x})|+|\widetilde{y}-y|\Big)$ (A.3)
$\hskip 28.45274pt\sup_{(x,y)\in{\mathcal{S}}^{train}}\sup_{(\widetilde{x},\widetilde{y})\in{\mathcal{S}}^{test}}\Big(|f_{\underline{\theta}_{*}}(x)|+|f_{\underline{\theta}_{*}}(\widetilde{x})|+|\widetilde{y}|+|y|\Big)$
By Lipschitz continuity of $f_{\underline{\theta}_{*}}$, there exists a constant $c_{0}:={\rm Lip}(f_{\underline{\theta}_{*}})$ such that
$x~|\displaystyle|f_{\underline{\theta}_{*}}(x)-f_{\underline{\theta}_{*}}(\widetilde{x})|\leq c_{0}$|x-$\widetilde${x}| (A.4)
and
$|f_{\underline{\theta}_{*}}(x)-f_{\underline{\theta}_{*}}(0)|\leq c_{0}|x|$ (A.5)
so that
$|f_{\underline{\theta}_{*}}(x)|\leq|f_{\underline{\theta}_{*}}(0)|+c_{0}|x|\,.$ (A.6)
Hence, the above is bounded by
$\leq$ $\sup_{(x,y)\in{\mathcal{S}}^{train}}\sup_{(\widetilde{x},\widetilde{y})\in{\mathcal{S}}^{test}}\Big(c_{0}|x-\widetilde{x}|+|\widetilde{y}-y|\Big)$ (A.7)
$\hskip 28.45274pt\sup_{(x,y)\in{\mathcal{S}}^{train}}\sup_{(\widetilde{x},\widetilde{y})\in{\mathcal{S}}^{test}}\Big(2|f_{\underline{\theta}_{*}}(0)|+c_{0}(|x|+|\widetilde{x}|)+|\widetilde{y}|+|y|\Big)$
$\leq$ $(1+c_{0})\sup_{(x,y)\in{\mathcal{S}}^{train}}\sup_{(\widetilde{x},\widetilde{y})\in{\mathcal{S}}^{test}}\Big(|x-\widetilde{x}|+|\widetilde{y}-y|\Big)$
$\hskip 28.45274pt2(1+|f_{\underline{\theta}_{*}}(0)|+c_{0})\sup_{(x,y)\in{\mathcal{S}}^{train}}\sup_{(\widetilde{x},\widetilde{y})\in{\mathcal{S}}^{test}}\Big(1+|x|+|\widetilde{x}|+|\widetilde{y}|+|y|\Big)$
$\leq$ $2\sqrt{2}(1+|f_{\underline{\theta}_{*}}(0)|+c_{0})^{2}(1+2R)\sup_{(x,y)\in{\mathcal{S}}^{train}}\sup_{(\widetilde{x},\widetilde{y})\in{\mathcal{S}}^{test}}\sqrt{|x-\widetilde{x}|^{2}+|\widetilde{y}-y|^{2}}$
$\leq$ $C_{0}(1+R)\;{\rm diam}({\mathcal{S}}^{train}\cup{\mathcal{S}}^{test})$
where $R$ is the radius of the smallest ball centered at the origin of ${\mathbb{R}}^{M_{0}}\times{\mathbb{R}}^{Q}$ that contains ${\mathcal{S}}^{train}\cup{\mathcal{S}}^{test}$.
We infer from the above steps that the constant $C_{0}$ is bounded by
$C_{0}\leq 8(1+|f_{\underline{\theta}_{*}}(0)|+c_{0})^{2}$ (A.8)
where $c_{0}$ is the Lipschitz constant of $f_{\underline{\theta}_{*}}$. Control of $C_{0}$ in (A.8) is obtained from Lemma A.1, below. ∎
Lemma A.1.
The Lipschitz constant $c_{0}$ of $f_{\underline{\theta}_{*}}$ is bounded by
$c_{0}\leq c_{\sigma}^{L}\prod_{\ell=1}^{L+1}\|W_{\ell}^{*}\|_{op}$ (A.9)
where $c_{\sigma}$ is the Lipschitz constant of $\sigma$, $\|\;\cdot\;\|_{op}$ is the matrix operator norm, and $(W_{\ell}^{*})_{\ell}$ are the minimizing weights for the training loss.
Moreover, recalling the bounds (1.2) on $\sigma$,
$|f_{\underline{\theta}_{*}}(0)|\leq C_{1}(a_{0},a_{1},\{\|W_{\ell}^{*}\|_{op},|b_{\ell}^{*}|\}_{\ell=1}^{L+1})$ (A.10)
is bounded by a constant depending on the operator norms of the trained weights, and the Euclidean norms of the trained biases.
Proof.
The map (1.1) in the $\ell$-th layer is Lipschitz continuous,
$|\sigma(W_{\ell}^{*}x_{1}^{(\ell-1)}+b_{\ell}^{*})-\sigma(W_{\ell}^{*}x_{2}^{(\ell-1)}+b_{\ell}^{*})|$ (A.11)
$\leq$ $c_{\sigma}|W_{\ell}^{*}x_{1}^{(\ell-1)}-W_{\ell}^{*}x_{2}^{(\ell-1)}|$
$\leq$ $c_{\sigma}\|W_{\ell}^{*}\|_{op}|x_{1}^{(\ell-1)}-x_{2}^{(\ell-1)}|$
with Lipschitz constant $c_{\sigma}\|W_{\ell}^{*}\|_{op}$. We also recall the basic fact that if a pair of functions $f_{1},f_{2}$ has Lipschitz constants $c_{1},c_{2}$ and the composition $f_{1}\circ f_{2}$ is well defined, then $f_{1}\circ f_{2}$ has a Lipschitz constant bounded by $c_{1}c_{2}$. By recursion in $\ell=1,\dots,L+1$, we arrive at the assertion.
Finally, recalling (1.2), we obtain
$|x^{(\ell)}|$ $\leq$ $|\sigma(W_{\ell}^{*}x^{(\ell-1)}+b_{\ell}^{*})|$ (A.12)
$\leq$ $a_{0}+a_{1}|W_{\ell}^{*}x^{(\ell-1)}+b_{\ell}^{*}|$
$\leq$ $a_{0}+a_{1}(\|W_{\ell}^{*}\|_{op}|x^{(\ell-1)}|+|b_{\ell}^{*}|)\,.$
Using $x^{(0)}=0$, recursion implies that (recalling that the output layer contains no activation function)
$|f_{\underline{\theta}_{*}}(0)|$ $=$ $|x^{(L+1)}|$ (A.13)
$\leq$ $\|W_{L+1}^{*}\|_{op}|x^{(L)}|+|b_{L+1}^{*}|$
$\leq$ $\|W_{L+1}^{*}\|_{op}(a_{0}+a_{1}(\|W_{L}^{*}\|_{op}|x^{(L-1)}|+|b_{L}^{*}|))+|b_{L+1}^{*}|$
$\leq$ $\cdots\;\leq\;C_{1}(a_{0},a_{1},\{\|W_{\ell}^{*}\|_{op},|b_{\ell}^{*}|\}_{\ell=1}^{L+1})$
is bounded by a constant depending on the operator norms of the trained weights, and the norms of the trained biases. ∎
Remark A.2.
The bounds proven in Lemma A.1 are consistent with the experimental observation that smaller weights increase generalization performance, through the decrease of the constants $c_{0}$ and $C_{1}$ in the current setting. In machine learning, training often aims to minimize both the loss and a term proportional to the $L^{2}$ norm of all parameters. This can be achieved through weight decay, where this term is added to the loss function. Additionally, gradient-based methods are known to introduce this term as an implicit regularizer in some cases, [GSGP25, Var23].
Remark A.3.
Generalization bounds involving the spectral norms of weights and Lipschitz constants of the activation have also been obtained in the context of probabilistic estimates, see for instance [BFT17, NK19, NBS18] and the references therein.
Appendix B Proof of Theorem 1.3
Let $g_{\underline{\theta}_{*}}$ be the map from the first hidden layer to the output layer for the trained network; hence, we have that
$f_{\underline{\theta}_{*}}(x)=g_{\underline{\theta}_{*}}(W_{1}^{*}x+b_{1}^{*})\,.$ (B.1)
It follows that
$|f_{\underline{\theta}_{*}}(\widetilde{x})-f_{\underline{\theta}_{*}}(x)|$ $\leq$ $|g_{\underline{\theta}_{*}}(W_{1}^{*}\widetilde{x}+b_{1}^{*})-g_{\underline{\theta}_{*}}(W_{1}^{*}x+b_{1}^{*})|$ (B.2)
$\leq$ ${\rm Lip}(g_{\underline{\theta}_{*}})|W_{1}^{*}(\widetilde{x}-x)|\,.$
We now consider each term of
${\mathcal{E}}_{\underline{\theta}_{*}}^{test}$ $=$ $\frac{1}{m}\sum_{j=1}^{m}|f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})-\widetilde{y}_{j}|^{2}.$ (B.3)
We find that for any arbitrary training data point $(x_{i}^{(0)},y_{i})\in{\mathcal{S}}^{train}$, we have
$|f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})-\widetilde{y}_{j}|^{2}$ (B.4)
$=$ $|f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})-\widetilde{y}_{j}-(f_{\underline{\theta}_{*}}(x_{i}^{(0)})-y_{i})+\underbrace{(f_{\underline{\theta}_{*}}(x_{i}^{(0)})-y_{i})}_{=:r(x_{i}^{(0)},y_{i})}|^{2}$
$\leq$ $\big(|f_{\underline{\theta}_{*}}(\widetilde{x}_{j}^{(0)})-f_{\underline{\theta}_{*}}(x_{i}^{(0)})|+|\widetilde{y}_{j}-y_{i}|+|r(x_{i}^{(0)},y_{i})|\big)^{2}$
$\leq$ $\Big({\rm Lip}(g_{\underline{\theta}_{*}})|W_{1}^{*}(\widetilde{x}_{j}^{(0)}-x_{i}^{(0)})|+|\widetilde{y}_{j}-y_{i}|+|r(x_{i}^{(0)},y_{i})|\Big)^{2}$
$\leq$ $(1+\delta)\Big({\rm Lip}(g_{\underline{\theta}_{*}})|W_{1}^{*}(\widetilde{x}_{j}^{(0)}-x_{i}^{(0)})|+|\widetilde{y}_{j}-y_{i}|\Big)^{2}+\left(1+\frac{1}{\delta}\right)|r(x_{i}^{(0)},y_{i})|^{2}$
$\leq$ $(1+\delta)C_{\underline{\theta}_{*}}^{2}\Big(|W_{1}^{*}(\widetilde{x}_{j}^{(0)}-x_{i}^{(0)})|+|\widetilde{y}_{j}-y_{i}|\Big)^{2}$
$\hskip 120.00018pt+\left(1+\frac{1}{\delta}\right)|r(x_{i}^{(0)},y_{i})|^{2}$
with
$C_{\underline{\theta}_{*}}:=\max\{1,{\rm Lip}(g_{\underline{\theta}_{*}})\}\,.$ (B.5)
Here, we used (B.2) to pass to the third line, and the fourth line follows from the Peter–Paul inequality (or Young’s inequality).
For each $j$, we now define $(x_{i(j)}^{(0)},y_{i(j)})\in{\mathcal{S}}^{train}$ as the training data point minimizing the first term in (B.4),
$|W_{1}^{*}(\widetilde{x}_{j}^{(0)}-x_{i(j)}^{(0)})|+|\widetilde{y}_{j}-y_{i(j)}|=\min_{i=1,\dots,n}\Big(|W_{1}^{*}(\widetilde{x}_{j}^{(0)}-x_{i}^{(0)})|+|\widetilde{y}_{j}-y_{i}|\Big)\,.$ (B.6)
Therefore, we obtain
(B.3) $\leq$ $(1+\delta)C_{\underline{\theta}_{*}}^{2}\frac{1}{m}\sum_{j=1}^{m}\Big(|W_{1}^{*}(\widetilde{x}_{j}^{(0)}-x_{i(j)}^{(0)})|+|\widetilde{y}_{j}-y_{i(j)}|\Big)^{2}$ (B.7)
$\hskip 56.9055pt+\left(1+\frac{1}{\delta}\right)r(x_{i(j)}^{(0)},y_{i(j)})^{2}$
$=$ $(1+\delta)C_{\underline{\theta}_{*}}^{2}d_{C}^{2}({\mathcal{S}}^{test}_{W_{1}^{*}}\,|\,{\mathcal{S}}^{train}_{W_{1}^{*}})+\left(1+\frac{1}{\delta}\right)\mathcal{E}_{\underline{\theta}_{*}}^{testNN}$
where $d_{C}({\mathcal{S}}^{test}_{W_{1}^{*}}\,|\,{\mathcal{S}}^{train}_{W_{1}^{*}})$ is the Chamfer pseudodistance between the point clouds ${\mathcal{S}}^{test}_{W_{1}^{*}}$ and ${\mathcal{S}}^{train}_{W_{1}^{*}}$, and
$\mathcal{E}_{\underline{\theta}_{*}}^{testNN}:=\frac{1}{m}\sum_{j=1}^{m}r(x_{i(j)}^{(0)},y_{i(j)})^{2}$ (B.8)
is the average training error for points in $\mathcal{S}^{train}$ which minimize the Chamfer distance to the test points. [^1] Here, $NN$ stands for nearest neighbor. Note that training points might contribute more than once to $\mathcal{E}_{\underline{\theta}_{*}}^{testNN}$. This proves the theorem. ∎
Appendix C Proof of Theorem 1.4
We first construct zero training loss minimizers for the $L^{2}$ loss function, using a similar argument as first presented in [CM]. Recalling the notations (1.28), zero loss is attained if and only if in the output layer,
$W_{L+1}X^{(L)}=Y-B_{L+1}\,$ (C.1)
where $Y$ is the matrix of reference output vectors $y_{i}\in{\mathbb{R}}^{Q}$. We choose
$b_{L+1}=-2\alpha(1,\dots,1)^{T}\;\;\;,\;\;\alpha=\max_{i,j}|(Y_{ij})_{-}|\,$, (C.2)
so that
$Y-B_{L+1}\;\;\;\in{\mathbb{R}}^{Q\times n}_{+}$ (C.3)
has non-negative matrix components. In particular, if all components of $Y$ are non-negative, $\alpha=0$.
In general, we may choose an arbitrary
$W_{L+1}\in{\mathbb{R}}^{Q\times M_{L}}_{+}\;\;\;,\;\;{\rm rank}(W_{L+1})=Q$ (C.4)
of full rank, so that $W_{L+1}W_{L+1}^{T}\in{\mathbb{R}}^{Q\times Q}$ is invertible, with non-negative matrix components. Then, denoting the generalized inverse of $W_{L+1}$ by
$W_{L+1}^{+}:=W_{L+1}^{T}(W_{L+1}W_{L+1}^{T})^{-1}\;\;\;\in\;{\mathbb{R}}^{M_{L}\times Q}_{+}$ (C.5)
we obtain that
$X^{(L)}=W_{L+1}^{+}(Y-B_{L+1})\;\;\;\in{\mathbb{R}}^{M_{L}\times n}_{+}$ (C.6)
satisfies (C.1), and is guaranteed to have non-negative matrix components.
In the case at hand, it is convenient to make the explicit choice
$W_{L+1}=[{\bf 1}_{Q\times Q}\;\;0_{Q\times(M_{L}-Q)}]$ (C.7)
with generalized inverse
$W_{L+1}^{+}=\left[\begin{array}[]{c}{\bf 1}_{Q\times Q}\\ 0_{(M_{L}-Q)\times Q}\end{array}\right]\,$ (C.10)
where $0_{p\times q}\in{\mathbb{R}}^{p\times q}$ consists of zeros. We then find for (C.6)
$X^{(L)}=W_{L+1}^{+}(Y-B_{L+1})=\left[\begin{array}[]{c}Y-B_{L+1}\\ 0_{(M_{L}-Q)\times n}\end{array}\right]\;\;\;\in{\mathbb{R}}^{M_{L}\times n}_{+}\,.$ (C.13)
In particular, $W_{L+1}$ and $W_{L+1}^{+}$ have operator norms
$\|W_{L+1}\|_{op}\;=\;\|W_{L+1}^{+}\|_{op}\;=\;1\,.$ (C.14)
The subsequent steps are the same for this specific choice of $W_{L+1}$, or for an arbitrary choice of the form (C.4).
In the next higher, $L$-th layer, we have
$X^{(L)}=\sigma(W_{L}X^{(L-1)}+B_{L})\;\;\;\in{\mathbb{R}}^{M_{L}\times n}_{+}$ (C.15)
and choose
[…]
$b_{L}$ $=$ $0\,$, (C.16)
and we let
$X^{(L-1)}=\left[\begin{array}[]{c}X^{(L)}\\ 0_{(M_{L-1}-M_{L})\times n}\end{array}\right]\;\;\;\in{\mathbb{R}}^{M_{L-1}\times n}_{+}\,.$ (C.19)
By recursion, we choose
[…]
$b_{\ell}$ $=$ $0\,$, (C.20)
for $\ell=2,\dots,L$, so that
$X^{(\ell)}=\left[\begin{array}[]{c}X^{(L)}\\ 0_{(M_{\ell}-M_{L})\times n}\end{array}\right]\;\;\;\in{\mathbb{R}}^{M_{\ell}\times n}_{+}\,.$ (C.23)
Hence, in the layer $\ell=1$, we arrive at
$X^{(1)}=\left[\begin{array}[]{c}X^{(L)}\\ 0_{(M_{1}-M_{L})\times n}\end{array}\right]\;\;\;\in{\mathbb{R}}^{M_{1}\times n}_{+}\,.$ (C.26)
and
$X^{(1)}=\sigma(W_{1}X^{(0)}+B_{1})\,.$ (C.27)
Here, we note that since $X_{0}\equiv X^{(0)}\in{\mathbb{R}}^{M_{0}\times n}$ has full rank $n$, it follows that the matrix $X_{0}^{T}X_{0}\in{\mathbb{R}}^{n\times n}$ is invertible, and we solve (C.27) with
$W_{1}=X^{(1)}X_{0}^{+}\;\;\;,\;\;b_{1}=0\,$, (C.28)
where
$X_{0}^{+}=(X_{0}^{T}X_{0})^{-1}X_{0}^{T}\,$ (C.29)
denotes the generalized inverse of $X_{0}$. Recalling (C.10), we have
$W_{1}$ $=$ $\left[\begin{array}[]{c}W_{L+1}^{+}(Y-B_{L+1})X_{0}^{+}\\ 0_{(M_{1}-M_{L})\times M_{0}}\end{array}\right]$ (C.32)
$=$ $\left[\begin{array}[]{c}(Y-B_{L+1})X_{0}^{+}\\ 0_{(M_{1}-Q)\times M_{0}}\end{array}\right]\;\;\;\in\;{\mathbb{R}}^{M_{1}\times M_{0}}$ (C.35)
by collecting rows of zeros. While $W_{1}$ has entries of indefinite sign, the product
$W_{1}X_{0}=\left[\begin{array}[]{c}Y-B_{L+1}\\ 0_{(M_{1}-Q)\times n}\end{array}\right]\;\;\;\in\;{\mathbb{R}}^{M_{1}\times n}_{+}$ (C.38)
has non-negative components. Therefore, $\sigma$ acts on the latter as the identity, and we indeed verify that (C.27) is satisfied with (C.26) and $b_{1}=0$.
Therefore, we arrive at the explicit zero loss minimizer $\underline{\theta}_{*}\in{\mathbb{R}}^{K}$ consisting of
$W_{1}^{*}$ $=$ $\left[\begin{array}[]{c}(Y-B_{L+1}^{*})X_{0}^{+}\\ 0_{(M_{1}-Q)\times M_{0}}\end{array}\right]$ (C.41)
$W_{\ell}^{*}$ $=$ $[{\bf 1}_{M_{\ell}\times M_{\ell}}\;\;0_{(M_{\ell-1}-M_{\ell})\times M_{L}}]\;\;\;,\;\;\ell=2,\dots,L$
$W_{L+1}^{*}$ $=$ $[{\bf 1}_{Q\times Q}\;\;0_{Q\times(M_{L}-Q)}]$
$b_{\ell}^{*}$ $=$ $0\;\;\;,\;\;\ell=1,\dots,L$
$b_{L+1}^{*}$ $=$ $-\alpha(1,\dots,1)^{T}\;\;\;,\;\;\alpha=\max_{i,j}|(Y_{ij})_{-}|\,.$ (C.42)
Then, similarly as in Lemma A.1, we find that the Lipschitz constant of $g_{\underline{\theta}_{*}}$ is bounded by
${\rm Lip}(g_{\underline{\theta}_{*}})\leq c_{\sigma}^{L}\prod_{\ell=2}^{L+1}\|W_{\ell}^{*}\|_{op}=1\,$, (C.43)
where we used (C.14), and the fact that the Lipschitz constant for ReLU is $c_{\sigma}=1$. Therefore, the constant in (1.21) satisfies
$C_{\underline{\theta}_{*}}$ $=$ $\max\{1,{\rm Lip}(g_{\underline{\theta}_{*}})\}$ (C.44)
$=$ $1\,.$
The generalization bound then follows from (1.22). This proves the claim. ∎
Acknowledgments: T.C. thanks Adam Klivans and Vardan Papyan for inspiring discussions, and gratefully acknowledges support by the NSF through the grant DMS-2009800, and the RTG Grant DMS-1840314. P.M.E. was supported by the NSF through T.C. with the grant DMS-2009800.